OpenStudy (anonymous):

Find the derivative of f(x)=x^3cosx

OpenStudy (anonymous):

use productr ule : f'(x) = x^3 * -sinx + cosx * 3x^2 = x^2(3cosx - xsinx)

OpenStudy (anonymous):

product rule is if f(x) = f(u) * f(v) f ' (x) = f'(u)* f(v) + f'(v) * f(u)

OpenStudy (anonymous):

thank you (: i was just making sure i was correct.

OpenStudy (anonymous):

ok no probs